# Curvature Mapping Method: Mapping Lorentz Force in Orion A

### v2

@@ 3.1 Comparison of Orientations
To evaluate the accuracy of the curvature mapping method, we compare the orientations of the pseudo Lorentz force and the projected Lorentz force. The offset angle between them in the x-y plane is:
We analyzed the offset angles in each slice along the z-axis for three MHD simulations (weak, middle, and strong magnetic field states) to simulate the variety of visual angles observed in actual observations (256256256 points per simulation, see Fig. 2). As shown in Fig. 2, 40%, 50%, 50% of the points in weak, middle, and strong magnetic states ( 0.3, 0.5, and 1.3, Zhao et al. 2024), exhibit parallel alignment (), and 50%, 60%, 60% of these simulations show close to parallel alignment (). This result demonstrates that the curvature mapping method effectively traces the orientation of the projected Lorentz force in the most cases.
3.2 Strength Modification of Projected Lorentz Force
The curvature mapping method estimates the 2D projected Lorentz force based on the POS magnetic field morphology and the total magnetic field strength. A modified factor from projected effect is introduced to improve the accuracy:
As shown in Fig. 3, the ratio from simulations follows a nearly normal distribution, with a standard deviation of 0.3 orders of magnitude (around two times). Gaussian fitting yields peak values of 1.1, 1.2, and 1.5 for weak, middle, and strong magnetic field states, respectively. Thus, ranges from 1.1 to 1.5.
Above all, the projected Lorentz force can be calculated by Curvature mapping method as:
where is approximately 1.1 to 1.5, is the scalar term of total magnetic field strength, and is the vector term, magnetic curvature derived from the magnetic field structure. Therefore, we use the OMC-1 magnetic field with a classical 2D ”hourglass” structure (Pattle et al., 2017) as an example to show the application of the curvature mapping method.
@@ 4 Mapping Lorentz Force in Orion A
By applying the Curvature Mapping Method to magnetic field observations of OMC-1 (Falgarone et al., 2008; Pattle et al., 2017), we constructed a Lorentz force map as vector fields within the molecular cloud. The vector form of the projected Lorentz force is determined using the Curvature Mapping Method, which relies on both scalar term (magnetic field strength) and vector terms (magnetic curvature):
where the orientation of the projected Lorentz force is aligned with the direction of the magnetic curvature, and its strength is predominantly influenced by the magnitude of the magnetic field strength and the magnetic curvature.
@@
4.1 Lorentz Force structure
The orientations and structure of Lorentz force can be traced by the orientation of magnetic curvature derived by magnetic field morphology. As Fig. 4 shows, the Lorentz force exhibits a structured, bipolar geometry in OMC-1 with an hourglass-shaped magnetic field. The observed Lorentz force distribution highlights localized regions of high curvature in the magnetic field, where the force vectors often diverge in opposing directions, forming a bipolar pattern. This structure suggests that the Lorentz force plays a critical role in regulating gas dynamics and counteracting gravitational collapse in these regions.
In the context of star formation, this interplay between Lorentz force and gravity is crucial. The Lorentz force in the hourglass-shaped magnetic field acts as a counterbalance to gravity, slowing down the collapse of gas and dust in molecular clouds (Galli & Shu, 1993; Allen et al., 2003; Qiu et al., 2014). This magnetic support is particularly important in regions of high curvature, where the magnetic field’s resistance to distortion is strongest. By buffering against gravitational collapse, the Lorentz force helps regulate the rate of star formation, preventing rapid and unregulated fragmentation of the cloud (Hacar et al., 2017).
@@ 4.1 Lorentz Force structure
4.2 Estimated magnetic field strength
Direct measurement of magnetic field strength in the interstellar medium (ISM) remains challenging. However, indirect methods, such as assuming force balance between magnetic and gravitational forces, provide viable estimates. As Fig. 4 shows, the Lorentz force in OMC-1, oriented along the filament body, supports against gravitational collapse (Hacar et al., 2017). By equating the Lorentz force to the self-gravity, the magnetic field strength can be estimated as:
where the is the volume density of gravity (, due to the projected Lorentz force as the volume density of magnetic force. We use a gravitational collapse region, OMC-1 (Hacar et al., 2017; Wu et al., 2024) as an example to estimate its magnetic field strength.
We apply this method to the star-forming region OMC-1, where Orion-KL and Orion-South (Fig. 4) are the dominant gravitational sources. Using column density maps (N(H2) 1023.4 cm-2, see Fig. 4), we derive clump masses M1(Orion-KL) and M2 (Orion-South), separated by L 0.2 pc. The gravity is roughly computed as .
Assuming force balance, the magnetic field strength is:
The magnetic curvature in OMC-1 follows a Gaussian distribution with a mean value of 1.6 pc-1 (Fig. 4). This yields an estimated field strength around 0.5 mG, consistent with Zeeman splitting measurements (-0.7 mG, Crutcher et al. 2010) and Davis-Chandrasekhar-Fermi (DCF) methods combined with autocorrelation function (ACF) analysis (0.6 mG; Wu et al. 2024). The magnetic field strength can be estimated by magnetic curvature from curvature mapping method. Of course, magnetic curvature would have more applications, including estimating magnetic field strength.
4.3 Mapping vector of Lorentz force in OMC-1
@@
If a reliable magnetic field strength distribution is available, the Lorentz force can be mapped in vector form (see Fig. 5). Using the ST method (Skalidis & Tassis, 2021; Skalidis et al., 2021), the magnetic field strength distribution is derived based on density, non-thermal velocity dispersion, and the angular dispersion of the magnetic field orientation (details are shown in Fig. D).
By combining the distribution maps of magnetic field strength (; see Fig. 8) and magnetic curvature (; see Fig. 4), the Lorentz force vectors can be calculated using the following expression:
@@
where the correction factor is taken as 1.5. The resulting Lorentz force vector map is displayed in Fig. 5.
The orientation of the Lorentz force is dictated by the magnetic curvature, as discussed in Sect. 4.1. Fig. 4 shows that the Lorentz force is strongest in high-density regions, while it diminishes in low-density areas. The Lorentz force relates to the H2 column density in OMC-1 , presented as a power law distribution ().
When weighted by column density, the probability density function (PDF) of the Lorentz force approximates a normal distribution. A Gaussian fit to the PDF reveals an average Lorentz force magnitude of approximately G2,cm-1.
Above, the Curvature Mapping Method enables the determination of both the orientation of the Lorentz force and the magnetic curvature from reliable magnetic field observations. One of the applications is estimating the magnetic field strength of molecular clouds using the strength of magnetic curvature. Furthermore, when the spatial distribution of the magnetic field strength is reliably obtained, the method allows for the construction of a detailed vector map of the projected Lorentz force in the interstellar medium. In the future, the Curvature mapping method will develop more applications to study the magnetic field of the interstellar medium and star formation.
@@ 5 Conclusion
We propose a new technique, the Curvature Mapping Method, to plot the vector map of the projected Lorentz force in the interstellar medium on the plane of the sky. The method is based on the idea that the Lorentz force is closely related to magnetic tension, which depends on the morphology and strength of the magnetic field:
@@
We test the accuracy of the Curvature Mapping Method to trace the projected Lorentz force using MHD simulations. In most cases (around 50-60), the orientation of Lorentz force measured by the Curvature Mapping Method is closely aligned with the projected component of the true Lorentz force. Its strength is also comparable to the real value of the 2D projected Lorentz force, with an error of approximately 100.3 ( two times).
@@
Based on credible magnetic field morphology, such as from dust polarization, the Curvature Mapping Method can determine the magnetic curvature and orientation of the POS Lorentz force.
Based on magnetic curvature from magnetic field morphology, the magnetic field strength can be estimated
If a credible magnetic field strength map is obtained, the vector map of the Lorentz force can be derived.
We measure the Lorentz force in OMC-1, which exhibits an hourglass-shaped magnetic field structure. The Lorentz force morphology presents a bipolar structure and supports gravity. The average magnetic curvature in OMC-1 is approximately 1.6 pc-1. Using the single magnetic field strength derived from Zeeman splitting ( -720G), the average strength of the Lorentz force is calculated as G2,cm-1. Using the magnetic curvature and the assumption of balance between Lorentz force and gravity, the magnetic field strength in OMC-1 is estimated as 0.5 mG, in agree with previous studies(Crutcher et al., 2010; Wu et al., 2024). As the column density increases, the Lorentz force grows larger (). The weighted average Lorentz force is approximately G2,cm-1.
By effectively using the information in the polarization map, the Curvature Mapping Method provides direct insights into the role of B-field in star-forming regions and can be widely applied. Our spatially-resolved approach takes full advantage of the information in the map, leading to direction vision into the role of the magnetic field. With this unprecedented amount of polarization measurements gathered by modern instruments, by combining our results with other methods with spatial information, such as gravitational acceleration mapping (He et al., 2023), direct insights into the role of the magnetic field in regulating star formation can be obtained.
Appendix A Data
A.1 Observations
@@
where polarized angle varies from -90∘ to 90∘. This is the IAU convention. One has to use =0.5 arctan(-U, Q) to convert Planck measurement to this IAU convention. The magnetic field (after that, B-field) orientation can be obtained by adding 90∘ to the polarized angle: = + 90∘.
@@
A.2 Simulations
The numerical simulation of molecular cloud selected in this work is identified within the three simulations applied by the constrained transport MHD option in Enzo (MHDCT) code (Collins et al. 2010; Burkhart et al. 2015). The simulation conducted in this study analyzed the impact of self-gravity and magnetic fields on supersonic turbulence in isothermal molecular clouds, using high-resolution simulations and adaptive mesh refinement techniques. which detail in shown in (Collins et al. 2012; Burkhart et al. 2015, 2020). The Enzo simulation with three dimensionless parameters provides us with massive information on physical processes:
where is the rms velocity fluctuation, cs is sonic speed, is mean density, and is the mean magnetic field strength. With the same sonic Mach number and viral parameter in initial conditions, the difference only exists in initial magnetic pressure with various B-field strength, The as 0.2, 2, and 20 present the strong B-field state, medium B-field, and weak B-field state. The clouds are in three different evolutionary stages started with self-gravity existing at 0.6( 0.76 Myr), where represents 1.1 Myr (Collins et al. 2012; Burkhart et al. 2015). Due to the short timescale of evolution, the gravitational collapse could barely affect the B- relation in the simulation The size of these molecular clouds is around 4.63 pc (2563 pixels), with one-pixel size of approximately 0.018 pc (Burkhart et al. 2015).
In this Enzo simulation, the thermal energy is located at the low state ( = 9). Their timescale of evolution stage is 0.6 , around 0.76 Myr, which the timescale starts from when interstellar media collapse. Due to the short evolutionary time ( 3.85 10-21 g cm-3, or 821 cm-3), the gravitational energy less affects the system in simulation. With the similar mean density in these simulations, the different (0.2, 2, 20) present the different magnetic field states, strong B-field, stronger B-field (weaker than strong B-field but stronger than weak B-field), and weak B-field. The Alfvén Mach number in various simulations can be calculated by the energy ratio between the total magnetic energy and kinetic energy (Zhao et al. 2024):
Appendix B Lorentz Force and Magnetic Tension in Enzo simulation
For this study, We use the simulation data ( = 20, t 0.76 Myr) on this work, whose the Alfvenic Mach number as 1.2 and the density range could be close to the actual molecular clouds (Zhao et al. 2024). This simulation setup is chosen such that the results resemble the structure of an actual star-forming molecular cloud, which is super-Alfvenic and close to the high-density molecular cloud (Zhao et al. 2024). The curvature mapping method can measure the Lorentz force in the pinch magnetic field, which mainly occurs in the high-density molecular cloud, and strong extra force, like self-gravity, distorts the local magnetic field. The Lorentz force (), magnetic tension (, and pressure are measured in 3D space as Eq. 3 and Eq. 4. With the density increasing, Lorentz force stays at a similar value with the magnetic tension and presure (See Fig. 6). In the order of magnitude, the Lorentz force is comparable to the magnetic tension and pressure.
We compare the Lorentz force, magnetic tension, and magnetic pressure in 3D space (Collins et al. 2012; Burkhart et al. 2015, 2020), where the similarity of two vectors, Lorentz force and another, is defined as:
The similarity is close to the half of the offset angle between two vectors in geometry (see Fig. 1) The similarity and offset angle in the 2D plane (POS) will be needed in actual observation. The offset angle in 3D space has a different probability density (see Ap. C). To accurately obtain the offset angle, the distribution of the offset angle is weighted by the probability density function (see Ap. C). The offset angle between magnetic tension and Lorentz force in the 2D plane (similar to the POS) is around 11∘, and they have similar sizes (see Fig. 7). The offset angle between magnetic pressure and Lorentz force in the 2D plane (similar to the POS) is around 19∘ (see Fig. 7), which are also close to parallel (offset angle 30∘). The magnetic tension and pressure resemble the Lorentz force in both magnitude and orientation.
Appendix C Remove the projection effect in probability density distributions
@@
Appendix D Estimate Distribution of Magnetic Field Strength
Accurately measuring the distribution of magnetic field strength on POS is challenging due to limitations in existing methods, such as Zeeman splitting (Crutcher et al. 2010) and the Davis-Chandrasekhar-Fermi (DCF) method (Davis 1951; Chandrasekhar & Fermi 1953; Crutcher et al. 2004). Zeeman splitting hardly maps magnetic field strength effectively which observations usually come from a single point source due to a low signal-to-noise ratio. The DCF method () relies on the assumption of equipartition between turbulent and magnetic energy and turbulence cause the distortion of straight magnetic field line, introducing errors from various observational uncertainties such as density () derived from continuum data(Zhao et al. 2025), velocity dispersion (), and magnetic field orientation angle dispersion () derived from polarization measurements. The DCF combined with the Autocorrelation Function (ACF,(Houde et al. 2009, 2016)) effectively avoids the overestimation of magnetic field distortion from turbulence but only provides one value of magnetic field strength in single clouds.
To measure the scale of POS Lorentz force, we choose the ST method (Skalidis & Tassis 2021; Arzoumanian et al. 2021) to estimate the plane distribution of magnetic field strength:
where the is the density of ISM, the as velocity dispersion relates to turbulence, is the magnetic field orientation angle dispersion which is affected by small-scale turbulence. The assumption of ST method (Skalidis & Tassis 2021; Arzoumanian et al. 2021) is similar to the DCF method (Davis 1951; Chandrasekhar & Fermi 1953; Crutcher et al. 2004). Different occur in the handle of angle dispersion , where that in ST method is and in DCF method is (DCF method : ). Regarding error propagation, the ST method could reduce the error from angle dispersion to magnetic field strength compared to the DCF method. Thus, we use the ST to estimate the plane distribution of magnetic field strength in OMC-1.
As Fig. 8 shows, we calculate the basic parameters, density , nor-thermal velocity dispersion , and magnetic field orientation angle dispersion in the same scale (, 0.07 pc). The volume density comes from the column density observed by Herschel (Poglitsch et al. 2010; Griffin et al. 2010; Roy et al. 2013; Polychroni et al. 2013) (see Ap. A), using the Constrained Diffusion Method (Li 2022) and its derivative method (Zhao Li +, in prep.) The angle dispersion is calculated by magnetic field orientation derived by 850m dust polarization (Pattle et al. 2017) and mapped by the sub-block of 33 beams (Zhao Li +, in prep.). The nor-thermal velocity dispersion comes from the velocity dispersion and kinetic temperature of NH3 observed by GBT (Friesen et al. 2017).

### v3

@@ 3.3.1 Comparison of Directions
We first evaluate the directional accuracy by calculating the offset angle between the estimated projected Lorentz force () and the projected Lorentz force () in the image plane.
Figure 4 displays the global distribution of offset angles for weak, middle, and strong magnetic field cases. Globally, approximately 50-60 of the pixels exhibit a ”close-to-parallel” alignment (). However, a critical trend is revealed when analyzing the error as a function of gas density (Fig. 4, bottom panels). In diffuse regions (volume density 102 cm-3), turbulence randomizes the magnetic structure, leading to large misalignments (30∘. In contrast, in high-density regions (), the alignment significantly improves, with the mean offset angle dropping to –. This density-dependent performance confirms that while the method may be limited in the diffuse ISM, it effectively and accurately traces the orientation of the Lorentz force in the dense molecular structures relevant to star formation.
3.3.2 Strength Modification and the Factor
Since the Curvature Mapping Method uses a geometric approximation for the force, a modification factor is introduced to account for the magnetic pressure force contribution and projection effects:
As shown in Fig. 5, the ratio follows a log-normal distribution with a dispersion of dex. Gaussian fitting yields mean values of , , and for magnetic, magnetic-kinetic, kinetic region , , and for strong, middle and weak magnetic states, respectively. Crucially, the bottom panels of Fig. 5 demonstrate that this ratio remains remarkably stable across the high-density regime. This stability justifies the use of a constant (ranging from 1.1 to 1.5) to recover the magnitude of the projected Lorentz force in dense cores.
In summary, the projected Lorentz force can be reliably mapped as:
where is the curvature vector derived from polarization maps. Given the method’s proven accuracy in dense gas, we proceed to apply it to the OMC-1 region, which features a classical 2D ”hourglass” magnetic field structure (Pattle et al., 2017).
@@ 4 Mapping Lorentz Force in Orion A
By applying the Curvature Mapping Method to magnetic field observations of OMC-1 (Falgarone et al., 2008; Pattle et al., 2017), we constructed a Lorentz force map as vector fields within the molecular cloud. The vector form of the projected Lorentz force is determined using the Curvature Mapping Method, which relies on both scalar term magnetic field strength and vector terms magnetic curvature:
where the orientation of the projected Lorentz force is aligned with the direction of the magnetic curvature, and its strength is predominantly influenced by the magnitude of the magnetic field strength and the magnetic curvature. We calculate the local curvature purely from the geometric morphology of the magnetic field lines. Specifically, we derive the orientation angle from the observed Stokes and parameters. We then apply Equations 5, 6, and 7 to compute the spatial derivatives of , yielding the projected curvature vector. This calculation is performed pixel-by-pixel for all regions satisfying our signal-to-noise criteria ( and ), without any prior assumptions about the magnetic field strength or Lorentz force
@@
4.1 Spatial Structure of Lorentz Force
The directions and structure of Lorentz force can be traced by the orientation of magnetic curvature derived by magnetic field morphology. As Fig. 6 shows, the Lorentz force exhibits a structured, bipolar geometry in OMC-1 with an hourglass-shaped magnetic field. The observed Lorentz force distribution highlights localized regions of high curvature in the magnetic field, where the force vectors often diverge in opposing directions, forming a bipolar pattern. This structure suggests that the Lorentz force plays a critical role in regulating gas dynamics and counteracting gravitational collapse in these regions.
In the context of star formation, this interplay between Lorentz force and gravity is crucial. The Lorentz force in the hourglass-shaped magnetic field acts as a counterbalance to gravity, slowing down the collapse of gas and dust in molecular clouds (Galli and Shu, 1993; Allen et al., 2003; Qiu et al., 2014). This magnetic support is particularly important in regions of high curvature, where the magnetic field’s resistance to distortion is strongest. By buffering against gravitational collapse, the Lorentz force helps regulate the rate of star formation, preventing rapid and unregulated fragmentation of the cloud (Hacar et al., 2017).
@@ 4.1 Spatial Structure of Lorentz Force
4.2 Lorentz Force vs. Gravity in OMC-1
4.2.1 Mapping Lorentz force in OMC-1
@@
If a reliable magnetic field strength distribution is available, the Lorentz force can be mapped in vector form (see Fig. 7). We employ the ST method (Skalidis and Tassis, 2021; Skalidis et al., 2021) to derive the scalar magnetic field strength map based on gas density, non-thermal velocity dispersion, and the angular dispersion of the magnetic field orientation (input parameter details are shown in Ap. C).
To derive the magnitude of Lorentz force, an independent measurement of the magnetic field strength () is required as an input. Therefore, we utilize the Structure Function (ST) method to derive the mean magnetic field strength (see Fig. 8). By combining the spatial distribution of magnetic field strength () and the vector map of magnetic curvature (; see Fig. 6), the Lorentz force vectors are calculated as:
@@
where the correction factor is taken as 1.5, consistent with the value derived for the weak-field case (see Sect.,3.3.2, Wu et al. 2024; Zhao et al. 2024). The resulting Lorentz force vector field is displayed in Fig. 7 (red arrows).
4.2.2 Spatially Resolved Gravity Vectors
To compare the Lorentz force and gravity, we calculated the gravitational potential () directly from the column density map and its gradient, gravitational acceleration. Following the 2D gravitational potential measurement (He et al., 2023), we solved the Poisson equation in Fourier space assuming a simplified geometry appropriate for the projected data:
where is the Fourier transform of the mass surface density (), is the wave vector, and represents an effective half-thickness (0.2 pc, filament width). This approach allows us to map the gravitational acceleration on POS:
The local volume density of the gravitational force is obtained by coupling the acceleration with the volume density (, derived in Ap.C):
where is volume density, detail shown in Fig. 8. The vector of gravitational force is plotted as blue vector in Fig. 7.
4.2.3 Magnetic Support Effectiveness
The spatially resolved maps of and allow for a direct, pixel-by-pixel evaluation of the interplay between magnetic fields and gravity. In contrast to the polarization orientation with limited dynamic range (), both the Lorentz force and gravitational force are polar vectors defined over the full range. This enables us to determine whether the magnetic force is supporting against ( close to 180∘) or assisting the gravitational collapse ( close to 0∘).
Fig. 7 presents the distribution of the alignment angle, , as a function of column density. A clear bimodal behavior is observed. In the diffuse envelope ( cm-2), the distribution is broad with a mean angle near , indicating that the magnetic force hardly affect gravity. In the dense filament ridge ( cm-2), the alignment transitions sharply to an anti-parallel configuration, with the angle approaching . The magnetic field is resisting gravitational collapse in the dense gas.
To quantify this interplay between gravity and magnetic force, we calculate the support effectiveness ratio, defined as the projection of the Lorentz force onto the direction opposing gravity, normalized by the gravitational force magnitude:
As shown in the Fig. 7, the support ratio is close to unity in the high-density ridge, where the projected component of the Lorentz force is comparable to the magnitude of gravity Considering the typical factor-of-two uncertainty in magnetic field strength estimation (which translates to a factor of uncertainty in force magnitude), the mean ratio in the ridge reaches . Spatially, the support effectiveness map (Fig. 7, middle panel) reveals that the ”backbone” of the Orion A filament (shown in red) experiences strong magnetic support (), while the surrounding regions (shown in blue) are dominated by gravitational infall. This suggests that while the region may be collapsing (Hacar et al., 2017), the magnetic force provides substantial support that likely regulates the fragmentation and collapse rate of the filament.
4.3 Constraints on Magnetic Field Strength
Direct measurement of magnetic field strength in the interstellar medium (ISM) remains challenging. However, the Curvature Mapping Method provides a way to estimate the equilibrium magnetic field strength () required to balance the local self-gravity. By equating the magnitude of the Lorentz force density () to the gravitational force density (), the magnetic field strength can be constrained as:
which yields:
This represents the theoretical magnetic field strength needed to support the cloud against gravitational collapse. For collapsing regions, this value serves as a physical reference for the field strength required to maintain hydrostatic equilibrium.
We apply this estimation to the OMC-1 region. Using the spatially resolved gravitational force map derived in Sect. 4.2.2 (see Fig. 7), the average gravitational force density in the dense ridge is approximately dyne cm-3. Combining this with the average magnetic curvature derived from the CMM ( pc-1, see Fig. 6) and adopting , Eq. 8 yields a required field strength of mG.
This estimated equilibrium value is remarkably consistent with independent measurements, including Zeeman splitting ( mG, Crutcher et al. 2010) and the Davis-Chandrasekhar-Fermi (DCF) method ( mG; Wu et al. 2024). This consistency suggests that the magnetic field in OMC-1 is strong enough to be comparable to gravity, yet the region remains dynamically active.
@@ 5 Conclusion
We propose a new technique, the Curvature Mapping Method (CMM), to obtain the vector map of the projected Lorentz force in the interstellar medium on the plane of the sky. The CMM effectively traces the restoring magnetic forces, provided that the magnetic field structure is shaped by the competition against gravity (or other external forces), rather than being in a force-free configuration. The method is based on the idea that the Lorentz force is closely related to magnetic tension, which depends on the morphology and strength of the magnetic field:
@@
We verified the accuracy of the method using 3D MHD simulations. While the method shows moderate accuracy in the diffuse ISM, it demonstrates remarkable reliability in star-forming regions. Specifically, in high-density gas ( cm-3), the derived force direction aligns with the true projected Lorentz force (offset ), and the magnitude is recovered within a factor of .
@@
Force Orientation: Based on credible magnetic field morphology (e.g., from dust polarization), the CMM can accurately determine the magnetic curvature and the orientation of the POS Lorentz force
Field Strength Constraint: Based on the magnetic curvature, the magnetic field strength required to balance gravity can be estimated, providing a theoretical upper limit for collapsing regions.
Vector Force Mapping: If a credible magnetic field strength map is available, the full vector map of the Lorentz force can be derived.
Applying the CMM to the OMC-1 region, combined with spatially resolved magnetic field strength (derived via the ST method) and self-gravity maps, we performed a pixel-by-pixel analysis of the dynamic interplay between magnetic fields and gravity. Our main findings are:
Vector Mapping: We successfully constructed the first 2D vector map of the Lorentz force in the Orion A filament.
Bimodal Behavior: We identified a clear transition in the magnetic force’s role. In the diffuse envelope, the magnetic force is statistically uncorrelated with gravity.
Magnetic Support: Along the dense filamentary ridge, the Lorentz force vectors are systematically anti-parallel to the gravitational acceleration vectors. The quantitative support ratio is found to be .
These results provide direct observational evidence that the magnetic field offers substantial support against gravitational collapse in the dense backbone of OMC-1, likely regulating the star formation rate. With modern polarization instruments (e.g., JCMT, ALMA), the Curvature Mapping Method offers a powerful new tool to visualize and quantify the dynamical role of magnetic fields in the interstellar medium.
Appendix A Observational Data
@@
where polarized angle varies from -90∘ to 90∘. This is the IAU convention. One has to use =0.5 arctan(-U, Q) to convert Planck measurement to this IAU convention (Planck Collaboration et al. 2020b). The magnetic field (after that, B-field) orientation can be obtained by adding 90∘ to the polarized angle: = + 90∘.
@@
Appendix B Correcting Projection Effects in Probability Density Distributions
@@
Appendix C Estimation of Magnetic Field Strength Distribution
Accurately measuring the distribution of magnetic field strength on POS is challenging due to limitations in existing methods, such as Zeeman splitting (Crutcher et al. 2010) and the Davis-Chandrasekhar-Fermi (DCF) method (Davis 1951; Chandrasekhar and Fermi 1953; Crutcher et al. 2004). Zeeman splitting hardly maps magnetic field strength effectively which observations usually come from a single point source due to a low signal-to-noise ratio. The DCF method () relies on the assumption of equipartition between turbulent and magnetic energy and turbulence cause the distortion of straight magnetic field line, introducing errors from various observational uncertainties such as density () derived from continuum data(Zhao et al. 2025), velocity dispersion (), and magnetic field orientation angle dispersion () derived from polarization measurements. The DCF combined with the Autocorrelation Function (ACF,(Houde et al. 2009, 2016)) effectively avoids the overestimation of magnetic field distortion from turbulence but only provides one value of magnetic field strength in single clouds.
To measure the scale of POS Lorentz force, we choose the ST method (Skalidis and Tassis 2021; 2021AA...647A..78A) to estimate the plane distribution of magnetic field strength:
where the is the density of ISM, the as velocity dispersion relates to turbulence, is the magnetic field orientation angle dispersion which is affected by small-scale turbulence. The assumption of ST method (Skalidis and Tassis 2021; 2021AA...647A..78A) is similar to the DCF method (Davis 1951; Chandrasekhar and Fermi 1953; Crutcher et al. 2004). Different occur in the handle of angle dispersion , where that in ST method is and in DCF method is (DCF method : ). Regarding error propagation, the ST method could reduce the error from angle dispersion to magnetic field strength compared to the DCF method. Thus, we use the ST to estimate the plane distribution of magnetic field strength in OMC-1.
As Fig. 8 shows, we calculate the basic parameters, density , non-thermal velocity dispersion , and magnetic field orientation angle dispersion in the same scale ( 40′′, 0.07 pc). The volume density comes from the column density observed by Herschel (Poglitsch et al. 2010; Griffin et al. 2010; Roy et al. 2013; Polychroni et al. 2013) (see Ap. A), using the Constrained Diffusion Method (Li 2022) and its derivative method (Zhao et al. 2025) The angle dispersion is calculated by magnetic field orientation derived by 850m dust polarization (Pattle et al. 2017) and mapped by the sub-block of 33 beams (Zhao et al. 2025).
The non-thermal velocity dispersion comes from the velocity dispersion and kinetic temperature of NH3 observed by GBT (Friesen et al. 2017).